2 x 13 x 4/15 - 5/6
Answer by stanbon(75887)

2 x 13 x 4/15 - 5/6
= 26 * (4/15) - (5/6)
= (104/15) - (5/6)
= (208/30) - (25/30)
= (183/30)
= 6.1
==============
Cheers,
Stan H.